OpenStudy (anonymous):

Simplify 3 sqrt 6 - 4 sqrt 6

OpenStudy (anonymous):

This is what we're dealing with correct? \[3\sqrt{6}-4\sqrt{6}\]

OpenStudy (anonymous):

yeah

OpenStudy (jdoe0001):

\(\bf 3{\color{brown}{ \sqrt{6}}}-4{\color{brown}{ \sqrt{6}}}\implies 3{\color{brown}{ \square }}-4{\color{brown}{ \square }}\implies ?\)

OpenStudy (anonymous):

3-4 = -1 -1 sqrt6 = -sqrt6

OpenStudy (anonymous):

\[3-4=-1\] \[-1\sqrt{6}=-\sqrt{6}\]

OpenStudy (anonymous):

@MarquisJohnson

OpenStudy (anonymous):

-11 sqrt 21 -11 sqrt 21 show all work

OpenStudy (anonymous):

Don't forget about the medal.

OpenStudy (anonymous):

Also, please post this on a new question not on the same question box.

OpenStudy (anonymous):

\[-11\sqrt{21}-11\sqrt{21}\]

OpenStudy (anonymous):

\[-11+-11 = -22\]

OpenStudy (anonymous):

The sqrt will stay the same. Proving that this answer will be \[-22\sqrt{21}\]